OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 5) with a slope of 2 in standard form.

OpenStudy (anonymous):

just like the last one \[y-y_1=m(x-x_1)\] with \[m=2, x_1=-1, y=5\]

OpenStudy (anonymous):

im bad at euations!!!!

OpenStudy (anonymous):

try it , write it down as i wrote it. make the replacement and see what you get then we can solve for y

OpenStudy (saifoo.khan):

yes, i can see bad at equations*. lol

OpenStudy (saifoo.khan):

No worries, practice makes a man perfect. ;)

OpenStudy (anonymous):

1. x = -1, 2. y = 5 3. m = 2

OpenStudy (anonymous):

So now write slope intercept form y = mx + b so we get 1. 5 = 2(-1) + b 2. 5 = -2 + b 3. 7 = b so your equation is what? y = ?x + ?